# SynSUM – Synthetic Benchmark with Structured and Unstructured Medical Records

### v1

@@ 2 Data generation
$\mathcal{P}_{\text{{joint}}}(\text{{asthma, smoking, ..., antibio, \# days}})=\mathcal{P}(\text{{asthma}})\mathcal{P}(\text{{smoking}})\mathcal{P}(\text{{COPD}}\mid\text{{smoking}})\\ \mathcal{P}(\text{{hayf}})\mathcal{P}(\text{{season}})\mathcal{P}(\text{{pneu}}\mid\text{{asthma, COPD, season}})\mathcal{P}(\text{{cold}}\mid\text{{season}})\\ \mathcal{P}(\text{{dysp}}\mid\text{{asthma, smoking, COPD, pneu, hayf}})\mathcal{P}(\text{{cough}}\mid\text{{asthma, smoking, pneu, cold}})\\ \mathcal{P}(\text{{pain}}\mid\text{{cough, pneu, COPD, cold}})\mathcal{P}(\text{{fever}}\mid\text{{pneu, cold}})\\ \mathcal{P}(\text{{nasal}}\mid\text{{cold, hayf}})\mathcal{P}(\text{{policy}})\mathcal{P}(\text{{self-empl}})\mathcal{P}(\text{{antibio}}\mid\text{{policy, dysp, cough, pain, fever}})\\ \mathcal{P}(\text{{\#days}}\mid\text{{antibio, dysp, cough, pain, fever, nasal, self-empl}})$ (1)
Our general methodology for generating the artificial patient records is shown in Figure 2. First, we sample a tabular patient record from an expert-defined Bayesian network. Then, we create a prompt containing the background information and symptoms available in the tabular patient record. We use this to prompt a large language model (GPT-4o) to generate a clinical note, as well as a more compact version of the same note. This process is repeated 10.000 times, leaving us with a dataset containing both the tabular patient record and two versions of the clinical note describing the patient encounter.
We now zoom in on the two major parts of this data generating process. First, Section 2.1 describes how we defined the generative process, including the Bayesian network modeling the structured tabular variables in the domain of respiratory disorders. Then, Section 2.2 dives into the clinical note generation by the large language model.
2.1 Modeling structured tabular variables with a Bayesian network
Causal structure
We asked an expert to define a Directed Acyclic Graph (DAG) which (partially) models the domain of respiratory diseases in primary care. In this DAG, a directed arrow between two variables models a causal relation between them. The full DAG is shown in the top portion of Figure 2.
Central to the model are the diagnoses of pneumonia and common cold (also known as upper respiratory tract infection), which may give rise to five symptoms. The expert also modeled some relevant underlying conditions which may render a patient more predisposed to certain diagnoses or symptoms. Based on the symptoms experienced by a patient, a primary care doctor decides whether to prescribe antibiotics or not. The presence and severity of the symptoms, as well as the prescription of antibiotics as a treatment, influence the total number of days that the patient eventually stays home as a result of illness (the outcome). Finally, there are some variables which are not clinical that exert an external influence on the diagnoses, the treatment and the outcome. Table 1 summarizes all variables and their meaning, as well as their possible values. While this model does not completely describe the real world, we do believe it to be sufficiently realistic for the purpose of generating an artificial dataset.
@@ 2.1 Modeling structured tabular variables with a Bayesian network
Season external influence Season of the year winter/summer
@@
Policy external influence Whether the clinician has higher or lower prior inclination to prescribe antibiotics. Can be influenced by many factors, such as local policy in their general practice, their own caution towards antibiotics or level of experience. high/low
Self-employed external influence Whether the patient is self-employed, rendering them less inclined to stay home from work for longer periods. yes/no
@@
Probability distribution
To define a data generating mechanism from which we can sample synthetic patients, we turn the DAG into a Bayesian network by defining a joint probability distribution. In a Bayesian network, this joint distribution factorizes into the product of conditional probability distributions for each variable, as shown in Equation (1). We parameterize these using four different approaches.
@@
For categorical variables with many parents, it becomes infeasible to manually fill in the CPT in a clinically meaningful way, because of the large number of possible combinations of parent values. This is the case for the symptoms dyspnea, cough, pain and nasal in our Bayesian network. To circumvent this problem, one can use a Noisy-OR distribution (Koller and Friedman, 2009). The Noisy-OR model is commonly used to define the distribution of a variable $Y$ which depends on a set of causes $\{X_{1},\dots,X_{k}\}$. It rests on the assumption that the combined influence of the possible causes $\{X_{1},\dots,X_{k}\}$ on $Y$ is a simple combination of the influence of each $X_{i}$ on $Y$ in isolation. This is a reasonable assumption to make in the case of symptoms with multiple possible causes (parents in the Bayesian network): a symptom arises in a patient if any of its possible causes succeeds in activating the symptom through its own independent mechanism. As shown in Equation (2), the parameterization of the noisy-OR distribution rests on choosing the parameters $p_{i}$, which is the probability that a possible cause $X_{i}$ activates symptom $Y$. As a special case, $p_{0}$, also known as the leak probability, is the probability that symptom $Y$ is activated as the result of another unmodeled cause (ouside of all $X_{i}$’s). Note that $x_{i}$ in the equation is $1$ when the cause $X_{i}$ is present in the patient, and 00 if not. Equations (3) through (6) define such a Noisy-OR distribution for the symptoms dyspnea, cough, pain and nasal. Note that the symptom fever is fully defined through a CPT, since the expert was able to provide intuition on all possible combinations of its two parent values, eliminating the need for a Noisy-OR distribution.
@@
The joint probability distribution from Equation (1) is now fully specified. We can use this Bayesian network to randomly sample the tabular portion of a patient record top-down, starting from the root variables without parents at the top and continuing further down. Each value is sampled conditionally on the variable’s parents’ values, using the conditional distributions we have defined. We repeat this process 10.000 times, leaving us with 10.000 artificial patient records consisting of 16 tabular features.
2.2 Generating unstructured text with a large language model
Starting from the tabular portion of the patient record, we aim to generate a text describing this fictional patient encounter. The scenario we try to mimic goes as follows. The patient goes to the primary care physician, telling them their symptoms and possible underlying conditions, along with additional context on the severity of these symptoms, when they started, among other details. The physician takes descriptive notes during this consultation, writing down the (recent) history prescribed by the patient. Then, based on the patient’s described complaints, they conduct a physical examination, writing down all findings. Both parts together then form the textual description of the patient encounter.
Our goal is to simulate this process by asking a large language model (LLM) to write a clinical note based on the information found in the tabular portion of the dataset. In this section, we describe how to build up the prompt which was used for this task. All generated prompts are generally structured like the example shown in Figure 2 (except for the special cases, see later). Only the background variables (asthma, smoking, COPD and hay fever) and symptoms (dyspnea, cough, pain, nasal and fever) may be directly mentioned in the prompt. The diagnoses pneumonia and common cold are not yet known to the clinician when they are noting down the patient history and physical examination, so these are not directly mentioned in the prompt. They can however influence the descriptors that are included next to the symptoms, as will be explained later. The treatment and outcome are left out of the prompt as well, just like the external influence variables, as all of these are not known at the time of writing the note or irrelevant to the simulated physician. Additional example prompts can be found in Appendix A.
@@ 2.2 Generating unstructured text with a large language model
The first block of information in the prompt concerns the symptoms experienced by the patient. We do not list the full set of symptoms exhaustively. Even if a patient might experience a certain symptom, there is a possibility that they do not mention it to the clinician, or that the clinician does not find it noteworthy to write down. On the other hand, if a patient does not experience a symptom, it is not very likely that they will mention this to the physician, and the physician might have no reason to ask for the symptom either. We therefore ask our expert to list the probability of mentioning the symptom in a clinical note when the symptom is positive and when it is negative. Of course, this would not generalize to all physicians, but it helps to bring some variety and realism in the notes we generate. The probabilities are as follows:
@@
For each symptom, we randomly sample whether it is to be mentioned in the prompt, conditional on its value. As can be seen in Figure 2, we explicitly tell the model what symptoms to mention and which to steer clear from. We randomly permute the ordering of the symptoms in each prompt.
@@
There are 3629 out of 10.000 patients where all symptoms in the tabular record are “no”, meaning the patient does not experience any respiratory symptoms. If we used the same prompt as before, this would result in an unrealistic clinical note, since the note would simply list all symptoms the patient does not have, without giving an actual reason for the patient’s visit. Furthermore, there would be little variation in these notes. An example is shown in Figure 7 in the Appendix. For these cases, it makes more sense to assume that the patient visits for a complaint unrelated to the respiratory domain, such as back pain, stomach issues, a skin rash, etc. To generate these special cases, we use a special prompt, telling the LLM the patient does not experience any of the 5 respiratory symptoms. When the patient has at least one underlying health condition (which is the case in 239 out of 3629 special cases), we add this to the prompt in the same way as before, like the example in Figure 4. If not (i.e., for the remaining 3390 out of 3629 special cases), we tell the LLM not to mention any of those health conditions either, see the prompt in Figure 8 in the Appendix. The latter prompt asks for three clinical notes at once, encouraging the LLM to be more creative and not repeat the same scenario every time, as well as being a little more cost-effective. This is possible because of the prompt being non-specific to any of these 3390 patients. We then randomly distribute all generated texts to each tabular patient record within this subset.
Compact version
Real clinical notes can be challenging, often containing abbreviations, shortcuts and denser sentence structure. To make the notes more challenging, we asked the LLM to create such a compact version of each note through an additional prompt, as can be seen in Figure 2. Our dataset contains both the original note and the compact version of the note.
@@
As a large language model, we opted for OpenAI’s GPT-4o model, using the version released in May 2024 (OpenAI, 2024). We set the temperature to 1.2 to encourage some more variation in the notes, while at the same time keeping them realistic. Before providing the case-specific prompt, we set the following system message: “You are a general practitioner, and need to summarize the patient encounter in a clinical note. Your notes are detailed and extensive.” We set the max_tokens parameter to 1000. All other parameters were kept as their default value. Generating all 10.000 notes and their compact version cost around $130$\$.
3 Symptom predictor baselines
To measure the amount of information encoded in the data and set a baseline for subsequent information extraction tasks, we run various prediction models on both the tabular and textual part of the dataset. These models are trained to predict each of the five symptoms: dyspnea, cough, pain, fever and nasal.
  • $\mathcal{P}(\text{{sympt}}\mid\text{{all}})$: Predict the symptom given all other tabular features as evidence. This set includes the background, diagnoses, external influence, treatment, outcome and other symptoms.
[…]
  • $\mathcal{P}(\text{{sympt}}\mid\text{{realistic}})$: Predict the symptom given all tabular features which would be available as evidence in a realistic setting. We do not expect policy, self-employed and #days to be recorded in any kind of realistic patient record, and therefore leave them out of this evidence set. As in the no-sympt setting, we assume that we have not extracted any tabular symptoms from the text yet. Therefore, this set includes the background, diagnoses, season and treatment variables.
Apart from the tabular-only baselines, we also train some baselines that get to see the text at the input. Our neural-text classifier takes only the text as an input (in the form of a pretrained clinical sentence embedding) and outputs the probability that a symptom is mentioned in the text. Finally, we extend this text-only baseline by concatenating a numerical representation of the tabular features to the text embedding at the input, forming the neural-text-tab baseline. Again, we do this for each of the three evidence settings outlined above. Note that this is the only model that combines both the background knowledge available in the tabular features with the unstructured text to extract, and it does so in a naive way. Future work will focus on improving the performance of this model by exploiting the relations between any of the tabular concepts, as envisioned in Figure 1.
dyspnea cough pain nasal fever
BN-tab
- all 0.7370 0.7816 0.2386 0.7146 0.4864
- no-sympt 0.7153 0.7776 0.1312 0.7146 0.4384
- realistic 0.6698 0.7763 0.0280 0.7146 0.3594
XGBoost-tab
- all 0.6639 0.7848 0.4070 0.7130 0.4111
- no-sympt 0.6612 0.7779 0.3638 0.7146 0.4015
- realistic 0.6626 0.7798 0.3698 0.7146 0.3951
neural-text
- normal 0.9660 0.9595 0.8415 0.9602 0.9074
neural-text-tab
- normal + all 0.9526 0.9481 0.8096 0.9598 0.9113
- normal + no-sympt 0.9592 0.9530 0.8078 0.9550 0.9094
- normal + realistic 0.9544 0.9543 0.8303 0.9575 0.9033
neural-text
- compact 0.9383 0.9480 0.7828 0.9583 0.8904
neural-text-tab
- compact + all 0.9535 0.9384 0.7675 0.9566 0.9096
- compact + no-sympt 0.9363 0.9240 0.7984 0.9638 0.8760
- compact + realistic 0.9442 0.9422 0.7880 0.9606 0.8926
Table 6: F1-score obtained over the test set for each of our baseline models. The results for the text classifiers trained over the normal vs. the compact version of the notes are grouped together for readability. We report results for the mean embedding type, while results for the other embedding types can be found in Table 7. We underline the best result obtained by the tabular-only models for each symptom, while the best overall result per symptom is in shown in bold.
dyspnea cough pain nasal fever
normal
- hist 0.9399 0.9699 0.8310 0.9538 0.9048
- phys 0.9035 0.7948 0.6871 0.9526 0.8466
- mean 0.9660 0.9595 0.8415 0.9602 0.9074
- concat 0.9635 0.9660 0.8313 0.9606 0.9086
compact
- hist 0.9353 0.9557 0.8008 0.9522 0.9148
- phys 0.8747 0.7655 0.6492 0.9544 0.8161
- mean 0.9383 0.9480 0.7828 0.9583 0.8904
- concat 0.9358 0.9443 0.7849 0.9598 0.8998
Table 7: Results for the neural-text baseline using different embedding types at the input. We report F1-score over the test set for both the normal and compact versions of the notes. The best results per symptoms and per note version is shown in bold. Figure 9 in the Appendix compares these results visually.
3.1 Models
BN-tab
We provide the causal structure in Figure 2 to the Bayesian network, and learn all parameters in the conditional probability tables (CPTs), Noisy-OR distributions, logistic regression model and Poisson regression model from the training data. In each case, we use maximum likelihood estimation to estimate the parameters and fill in a CPT. Where we don’t directly learn a CPT (for the variables dyspnea, cough, pain, nasal, antibiotics and #days), we evaluate the learned distribution for each combination of child and parent values to obtain a CPT. For more details, we refer to Appendix B.1. We then use variable elimination over the full joint distribution to evaluate the capability of the learned Bayesian network to predict each of the symptoms, taking different variables as evidence according to the 3 settings described earlier (all, no-sympt and realistic).
XGBoost-tab
We train an XGBoost classifier for each symptom in combination with each of the 3 settings, meaning each classifier has a different set of tabular features at the input. We optimize the hyperparameters separately for each combination (15 in total) using 5-fold cross-validation. For more details, we refer to Appendix B.2.
Neural-text
We train a neural classifier that takes only the text as an input and is trained to predict the probability a symptom is mentioned. We train separate classifiers for each symptom. We first split the text into sentences, and transform these into an embedding using the pretrained clinical representation model BioLORD (Remy, Demuynck, and Demeester, 2024). We explore 4 settings for turning these sentence embeddings into a single note embedding:
The note embedding is then fed into a multi-layer perceptron with one hidden layer, followed by a Sigmoid activation for the symptoms dyspnea, cough, pain and nasal, and a Softmax activation with 3 outputs heads for the symptom $fever$. We optimized the parameters of each model (i.e. each combination of symptom and embedding type) using the binary or multiclass cross-entropy objective over the symptom labels. For more details, we refer to Appendix B.3.
Neural-text-tab
We extend the neural-text baseline by concatenating the mean text embeddings with the tabular variables at the input of each neural classifier. All categorical tabular variables were first transformed into a one-hot encoding, while the variable #days was preprocessed using standard scaling. We used the same architecture as the neural-text baseline (only changing the dimension of the input layer), and again trained separate classifiers for each symptom in combination with each evidence setting (all, no-sympt and realistic).
3.2 Results
We randomly split the dataset into a train and test set, using an 8000/2000 split. We use cross-validation on the train set to tune any hyperparameters, and report the final F1-score over the test set after training. We classify a symptom as positive if the predicted probability is larger than $0.5$. Since fever has three possible categories, the class with the highest predicted probability is chosen. In that case, we report the macro F1-score over all three categories.
Table 6 compares the results obtained over all baselines. The tabular-only baselines (BN-tab and XGBoost-tab) perform consistently worse than the baselines that include text (neural-text and neural-text-tab). The evidence setting where all other features are included as evidence usually performs best for the tabular-only baselines.
The neural-text-tab baseline does not perform better than the neural-text baseline when the normal notes are used. While there is little room for improvement in the dyspnea, cough and nasal classifiers, the symptoms pain and fever are harder to predict. We also note a consistent gap in performance between the normal notes and the compact notes, which can be attributed to the higher complexity of the compact notes. In that case, the neural-text-tab classifier manages to marginally improve over the neural-text classifier by including the tabular features.
Table 7 further breaks down the results for the neural-text classifier over the different embedding types. There is a significant gap in performance between the hist and phys settings for the symptoms cough, pain and fever. This makes sense, as the “history” section of the note outlines the symptoms experienced by the patient more clearly. The performance difference between the mean and concat settings is usually small, with concat slightly outperforming mean. While the score for hist comes close to those for mean and concat, the latter usually still outperform the former for the normal notes, showing that there is some complementary information in the “history” and “physical examination” portions. For the advanced notes, including the “physical examination” portion only seems to confuse the model, since the hist setting usually outperforms mean and concat there.
4 Discussion
Symptom extraction
Our analysis of simple tabular and textual baseline models revealed that the symptoms pain and fever are hardest to predict in both the tabular-only and the text-only setting. Combining both settings, i.e. integrating tabular background features in the extraction of the concepts from the text, linking them through domain knowledge, may have potential for improved information extraction. Future work will focus on realizing this hybrid approach to improve upon the baseline results presented in Table 6. The naive way of including tabular features by simply concatenating them to the text embedding at the input of the neural classifier already improves extraction performance on the more difficult compact version of the notes. This indeed suggests that symptoms that are mentioned less explicitly in the text might be recovered by including background tabular features in the prediction.
Limitations
While the dataset we constructed is meant to have realistic properties, we also intentionally simplify reality to make the design and generation process feasible. The dataset is purely meant as a research benchmark where the ground truth relations are known, and results obtained on it are not meant to transfer to real clinical notes or datasets. We therefore advise strongly against using the dataset for training prediction models which will be deployed in real settings.
Potential uses
The dataset is primarily designed to facilitate research on clinical information extraction in the presence of tabular background variables. Future work will focus on realizing the idea presented in Figure 1, where the tabular features aid in more accurately extracting concepts from the text by linking them through domain knowledge. Apart from this, we also foresee multiple secondary uses of the dataset. First, the dataset could facilitate research on the automation of clinical reasoning over tabular data and text, following the example of (Rabaey et al., 2024). Second, it could be used to benchmark causal effect estimation methods in the presence of textual confounders, similar to Veitch, Sridhar, and Blei (2020), thanks to the purposeful inclusion of both a treatment and outcome variable in our dataset. Third, there has been increasing interest in clinical synthetic data (Hernandez et al., 2022), where a set of patient characteristics is turned into a synthetic version that is meant to protect the privacy of individuals in the original dataset. Our dataset could serve as a benchmark for comparing synthetic data generation methods that jointly generate tabular variables and text (Lee, 2018; Ceritli et al., 2023; Guan et al., 2019). In short, any area of research focusing on the intersection of tabular data and text in healthcare can potentially benefit from our proposed benchmark.
Acknowledgments
Paloma Rabaey and Henri Arno’s research is funded by the Research Foundation Flanders (FWO Vlaanderen). This research also received funding from the Flemish government under the “Onderzoeksprogramma Artificiële Intelligentie (AI) Vlaanderen” programme.
Appendix
Appendix A Additional example prompts
Figures 5 and 6 show two additional example prompts. Figure 7 shows what would happen if we prompted the unrelated cases (where no symptoms are present in the patient) using our normal strategy. Figure 8 shows the prompt we used instead, in the case where the patient experiences no respiratory symptoms and does not have any underlying respiratory conditions either.

### v2

@@ Appendix B Bayesian network
$\mathcal{P}_{\text{{joint}}}(\text{{asthma, smoking, ..., antibio, \# days}})=\mathcal{P}(\text{{asthma}})\mathcal{P}(\text{{smoking}})\mathcal{P}(\text{{COPD}}\mid\text{{smoking}})\\ \mathcal{P}(\text{{hayf}})\mathcal{P}(\text{{season}})\mathcal{P}(\text{{pneu}}\mid\text{{asthma, COPD, season}})\mathcal{P}(\text{{cold}}\mid\text{{season}})\\ \mathcal{P}(\text{{dysp}}\mid\text{{asthma, smoking, COPD, pneu, hayf}})\mathcal{P}(\text{{cough}}\mid\text{{asthma, smoking, COPD, pneu, cold}})\\ \mathcal{P}(\text{{pain}}\mid\text{{cough, pneu, COPD, cold}})\mathcal{P}(\text{{fever}}\mid\text{{pneu, cold}})\\ \mathcal{P}(\text{{nasal}}\mid\text{{cold, hayf}})\mathcal{P}(\text{{policy}})\mathcal{P}(\text{{self-empl}})\mathcal{P}(\text{{antibio}}\mid\text{{policy, dysp, cough, pain, fever}})\\ \mathcal{P}(\text{{\#days}}\mid\text{{antibio, dysp, cough, pain, fever, nasal, self-empl}})$ (1)
@@
Season non-clinical Season of the year winter/summer
@@
Policy non-clinical Whether the clinician has higher or lower prior inclination to prescribe antibiotics. Can be influenced by many factors, such as local policy in their general practice, their own caution towards antibiotics or level of experience. high/low
Self-employed non-clinical Whether the patient is self-employed, rendering them less inclined to stay home from work for longer periods. yes/no
@@
To define a data generating mechanism from which we can sample synthetic patients, we turn the DAG from Figure 1 into a Bayesian network by defining a joint probability distribution. In a Bayesian network, this joint distribution factorizes into the product of conditional probability distributions for each variable, as shown in Equation (1). We parameterize these using four different approaches.
Conditional probability table
@@
Noisy-OR distribution
For categorical variables with many parents, it becomes infeasible to manually fill in the CPT in a clinically meaningful way, because of the large number of possible combinations of parent values. This is the case for the symptoms dyspnea, cough, pain and nasal in our Bayesian network. To circumvent this problem, we define a Noisy-OR distribution (Koller and Friedman 2009). The Noisy-OR model is commonly used to define the distribution of a variable $Y$ which depends on a set of causes $\{X_{1},\dots,X_{k}\}$. It rests on the assumption that the combined influence of the possible causes $\{X_{1},\dots,X_{k}\}$ on $Y$ is a simple combination of the influence of each $X_{i}$ on $Y$ in isolation. This is a reasonable assumption to make in the case of symptoms with multiple possible causes (parents in the Bayesian network): a symptom arises in a patient if any of its possible causes succeeds in activating the symptom through its own independent mechanism. As shown in Equation (2), the parameterization of the noisy-OR distribution rests on choosing the parameters $p_{i}$, which is the probability that a possible cause $X_{i}$ activates symptom $Y$. As a special case, $p_{0}$, also known as the leak probability, is the probability that symptom $Y$ is activated as the result of another unmodeled cause (ouside of all $X_{i}$’s). Note that $x_{i}$ in the equation is $1$ when the cause $X_{i}$ is present in the patient, and 00 if not. Equations (3) through (6) define such a Noisy-OR distribution for the symptoms dyspnea, cough, pain and nasal. Note that the symptom fever is fully defined through a CPT, since the expert was able to provide intuition on all possible combinations of its two parent values, eliminating the need for a Noisy-OR distribution.
@@
The joint probability distribution from Equation (1) is now fully specified. We can use this Bayesian network to randomly sample the tabular portion of a patient record top-down, starting from the root variables without parents at the top and continuing further down. Each value is sampled conditionally on the variable’s parents’ values, using the conditional distributions we have defined. We repeat this process $10{,}000$ times, leaving us with $10{,}000$ artificial patient records consisting of 16 tabular features.
Appendix C Prompting the large language model
Starting from the tabular portion of the patient record, we aim to generate a text describing this fictional patient encounter. The scenario we simulate artificially is as follows. The patient goes to the primary care physician, telling them their symptoms and possible underlying conditions, along with additional context on the severity of these symptoms, when they started, among other details. The physician takes descriptive notes during this consultation, writing down the (recent) history prescribed by the patient. Then, based on the patient’s described complaints, they conduct a physical examination, writing down all findings. Both parts together then form the textual description of the patient encounter.
@@ Appendix C Prompting the large language model
The first block of information in the prompt concerns the symptoms experienced by the patient. We do not list the full set of symptoms exhaustively. Even if a patient might experience a certain symptom, there is a possibility that they do not mention it to the clinician, or that the clinician does not find it noteworthy to write down. On the other hand, if a patient does not experience a symptom, it is not very likely that they will mention this to the physician, and the physician does not always have a reason to ask for the symptom either. We therefore ask our expert to list the probability of mentioning the symptom in a clinical note when the symptom is positive and when it is negative. Of course, this would not generalize to all physicians, but it helps to bring some variety and realism in the notes we generate. The probabilities are as follows:
@@
For each symptom, we sample whether it is to be mentioned in the prompt, conditional on its value, according to the probabilities stated above. As can be seen in Figure 1, we explicitly tell the model what symptoms to mention and which to steer clear from. We randomly permute the ordering of the symptoms in each prompt.
@@
Figure 4: Example of what would happen if we simply extended the general prompt to the 3629 cases where the patient does not experience any respiratory symptoms. There would be little variation in these notes, as these patients seem to visit the doctor’s office for no reason. It is more realistic to encourage the LLM to generate a note which describes a patient visiting for a non-respiratory complaint, like the prompts shown in Figures 5 and 6.
There are 3629 out of 10,000 patients where all symptoms in the tabular record are “no”, meaning the patient does not experience any respiratory symptoms. If we used the same prompt as before, this would result in an unrealistic clinical note, since the note would simply list all symptoms the patient does not have, without giving an actual reason for the patient’s visit. Furthermore, there would be little variation in these notes. An example is shown in Figure 4. For these cases, it makes more sense to assume that the patient visits for a complaint unrelated to the respiratory domain, such as back pain, stomach issues, a skin rash, etc. To generate these special cases, we use a special prompt, telling the LLM the patient does not experience any of the 5 respiratory symptoms. When the patient has at least one underlying health condition (which is the case in 239 out of 3629 special cases), we add this to the prompt in the same way as before, like the example in Figure 5. If not (i.e., for the remaining 3390 out of 3629 special cases), we tell the LLM not to mention any of those health conditions either, see the prompt in Figure 6. The latter prompt asks for three clinical notes at once, encouraging the LLM to be more creative and not repeat the same scenario every time, as well as being a little more cost-effective. This is possible because of the prompt being non-specific to any of these 3390 patients. We then randomly distribute all generated texts to each tabular patient record within this subset.
@@
As a large language model, we opted for OpenAI’s GPT-4o model, using the version released in May 2024 (OpenAI 2024). We set the temperature to 1.2 to encourage some more variation in the notes, while at the same time keeping them realistic. Before providing the case-specific prompt, we set the following system message: “You are a general practitioner, and need to summarize the patient encounter in a clinical note. Your notes are detailed and extensive.” We set the max_tokens parameter to 1000. All other parameters were kept as their default value. Generating all 10,000 notes and their compact version cost around $130$\$.
Appendix D Additional example prompts
Figures 7 and 8 show two additional example prompts.